# Born in the Dark: The Catastrophic Collapse of Fuzzy Dark Matter Solitons as the Origin of Little Red Dots

### v1

@@ III.2 Dust Attenuation and Broad Line Visibility
While electron scattering suppresses X-rays, dust provides the dominant opacity for UV/optical photons. We treat $A_{V}$ as an order-of-magnitude indicator of heavy reprocessing.
[…]
@@ III.3 Thermodynamic Equilibrium
Assuming the obscured luminosity is re-radiated in the IR, a characteristic effective temperature follows from the Stefan–Boltzmann law. For $L_{\rm bol}\sim 10^{46}\,{\rm erg\,s^{-1}}$ and a photospheric scale $R_{\rm ph}\sim{\cal O}(r_{c})$, we obtain $T_{\rm eff}\approx 150$$200\,{\rm K}$, implying a rest-frame mid-IR bump (peak near $\sim 20\,\mu{\rm m}$). The very red JWST colors at shorter wavelengths likely require an additional hotter dust component (near the sublimation radius) and/or anisotropic leakage/scattering through a non-uniform cocoon; quantifying the detailed SED requires radiative-transfer/radiation-hydrodynamic modeling.
@@ IV.1 Setup and Initial Conditions
We solve the Schrödinger–Poisson system using a pseudo-spectral code on a $N=512^{3}$ grid with periodic boundary conditions (see Schive et al., 2014a). Our aim is not to claim the discovery of soliton formation, but to demonstrate that compact cores with LRD-like scales form robustly in this setup. To model the formation of the potential well, we initialize the wavefunction with eight identical ground-state solitons with zero initial phases (free-fall merger). While we use zero phase for demonstration, previous studies have shown soliton formation is robust against random phase fluctuations (Schive et al., 2014b). The code is GPU-accelerated (CUDA) and was executed on an NVIDIA RTX 3080, requiring approximately one day of runtime.
@@ V Simulation Results
Figure 3: Validation of a compact soliton core in an idealized merger. Radial density profile of the simulated core (red points) plotted against the analytic soliton solution (black line). The inner region is well described by the soliton profile.
@@ VI Discussion
VI.1 The Tension with Lyman-$\alpha$ Forest
We acknowledge that $m_{22}\approx 2.0$ is in tension with some Lyman-$\alpha$ constraints ($m_{22}\gtrsim 20$). However, the LRD size-based argument is a direct constraint on the compact gravitational potential shape. Under the working assumptions adopted here ($M_{s}\sim 10^{8}$$10^{9}\,M_{\odot}$ and $r_{e}\sim\mathcal{O}(r_{c})$), the observed compact sizes favor $m_{22}\sim{\rm few}$. If Lyman-$\alpha$ forest analyses robustly require $m_{22}\gtrsim\mathcal{O}(10)$, possible resolutions include a sub-dominant FDM fraction or revisions to IGM thermal history assumptions.
@@ VI.2 Observational Signatures
  • Inverse Size–Mass Relation: For soliton cores at fixed $m_{22}$, the soliton scaling implies $r_{c}\propto M_{c}^{-1}$. Figure 4 serves as an illustrative mapping assuming $M_{\rm BH}$ tracks $M_{c}$. Establishing this as a discriminator will require joint constraints on host masses.
@@ VI.3 Demographics and Duty Cycle
A key question is whether the transient obscured phase implied by the “Opacity Crisis” can match the observed abundance of LRDs, $n_{\rm LRD}\sim 10^{-5}\,{\rm comoving\ Mpc^{-3}}$ (here cMpc denotes comoving Mpc). For massive halos at $z\sim 5$77, a representative number density is $n_{\rm halo}\sim 10^{-4}\,{\rm cMpc^{-3}}$ (e.g., Kokorev et al., 2024), suggesting an instantaneous occupancy fraction of order $f_{\rm occ}\equiv n_{\rm LRD}/n_{\rm halo}\sim 0.1$.
More generally, the occupancy can be written as a time-averaged duty cycle,
@@
where $P_{\rm trig}$ is the probability that a given halo experiences an LRD-triggering event within the relevant redshift window, $N_{\rm evt}$ is the number of such episodes per halo, $t_{\rm evt}$ is the observable duration of each episode, and $t_{H}$ is the Hubble time at that epoch. While the free-fall time of the dense core can be short ($t_{\rm ff}\sim 10^{5}\,{\rm yr}$), feedback, clumpy obscuration, and/or repeated episodes can extend the effective observable duration to $t_{\rm evt}\sim 10^{6}$$10^{7}\,{\rm yr}$. Matching $f_{\rm occ}\sim 0.1$ then requires $P_{\rm trig}N_{\rm evt}$ to be of order unity to several, motivating multiple merger-driven or inflow-driven obscured episodes during early assembly.
[…]
@@ VII Conclusion
We have presented a compact model in which LRD-like systems arise during a transient, heavily obscured phase inside FDM soliton cores. Combining size scaling with opacity constraints suggests $m_{22}\sim{\rm few}$. A conservative column estimate, including sensitivity analysis for baryon loading, indicates that reaching $N_{\rm H}\gtrsim 10^{24}\,{\rm cm^{-2}}$ typically implies radiative losses faster than dynamical support (or radiation-pressure dominance), disfavoring a long-lived static hot atmosphere and motivating rapid inflow or radiation-pressure-driven evolution.

### v2

@@ III.2 Dust Attenuation and Broad Line Visibility
Recent analyses suggest that LRDs are best described as young SMBHs embedded in compact, physically thick obscurers (“dense ionized cocoons” and/or heavy dust shrouds) with column densities $N_{\rm H}\gtrsim 10^{24}\,{\rm cm^{-2}}$ (e.g., Rusakov et al., 2026). Our FDM soliton framework provides a natural origin for such a geometry: unlike a rotationally supported thin disk, the soliton produces a quasi-spherical potential well on the characteristic scale $r_{c}\sim 50$100100 pc, so gas inflowing into the core can initially fill a large fraction of the volume.
A simple parameterized cocoon geometry.
We parameterize the obscurer as extending from an inner dust sublimation scale $r_{\rm in}$ to an observable effective radius
$r_{\rm e}\simeq\xi\,r_{c},$ (16)
with a line-of-sight covering factor $f_{\rm cov}\gtrsim 0.5$. A convenient phenomenological description is a biconical polar opening with half-opening angle $\theta$, for which $f_{\rm cov}\simeq\cos\theta$. In this picture, broad lines remain visible either through low-column sightlines (gaps) or via scattering into the line of sight, while the high covering factor naturally accounts for the high detection rate of obscured LRDs (e.g., Greene et al., 2024).
Granularity and clumpiness in the soliton core.
A specific ingredient of the FDM core is its intrinsic time-dependent granularity arising from wave interference. Order-unity density fluctuations (“granules”; $\delta\rho/\rho\sim 1$) produce persistent gravitational stirring on timescales comparable to the soliton dynamical/oscillation time. These perturbations can promote and help maintain a clumpy cocoon/torus morphology, providing a complementary route to clumpiness beyond magnetic or radiation-driven instabilities.
How cocoon observables can constrain the FDM scenario.
In this framework the “Opacity Crisis” describes the global collapse/evolution of a (possibly clumpy) atmosphere once $t_{\rm rad}\lesssim t_{\rm dyn}$. Over time, feedback may carve polar channels and reduce $f_{\rm cov}$, but the deeper potential of a soliton can qualitatively help retain a high-column equatorial component longer than a shallow, baryon-dominated halo. Observables such as the broad-line visibility fraction, polarization, and inferred covering factors/opening angles (and their evolution with mass and redshift) therefore provide potential empirical constraints on the depth of the central potential and, by extension, the allowed range of $m_{22}$ (see also Delvecchio et al., 2025).
@@ III.3 Thermodynamic Equilibrium
Assuming the obscured luminosity is re-radiated in the IR, a characteristic effective temperature follows from the Stefan–Boltzmann law. For $L_{\rm bol}\sim 10^{46}\,{\rm erg\,s^{-1}}$ and a photospheric scale $R_{\rm ph}\sim{\cal O}(r_{c})$, we obtain $T_{\rm eff}\approx 150$$200\,{\rm K}$, implying a rest-frame mid-IR bump (peak near $\sim 20\,\mu{\rm m}$). The very red JWST colors at shorter wavelengths likely require an additional hotter dust component (near the sublimation radius) and/or anisotropic leakage/scattering through a non-uniform cocoon; quantifying the detailed SED requires radiative-transfer/radiation-hydrodynamic modeling. In this context the soliton potential primarily sets the outer scale and confinement (through $r_{c}$ and $r_{\rm e}$); it does not by itself determine the dust microphysics at $r_{\rm in}$, so the need for a hot dust component and its dependence on halo/core properties ultimately requires dedicated radiation-hydrodynamic simulations.
@@ IV.1 Setup and Initial Conditions
We solve the Schrödinger–Poisson system using a pseudo-spectral code on a $N=512^{3}$ grid with periodic boundary conditions (see Schive et al., 2014a). Our aim is not to claim the discovery of soliton formation, but to demonstrate that compact cores with LRD-like scales form robustly in this setup. To model the formation of the potential well, we initialize the wavefunction with eight identical ground-state solitons with zero initial phases (free-fall merger). While we use zero phase for demonstration, previous studies have shown soliton formation is robust against random phase fluctuations (Schive et al., 2014b). The simulations were executed on an NVIDIA RTX 3080, requiring approximately one day of runtime.
@@ V Simulation Results
Figure 3: Validation of a compact soliton core in an idealized merger. Radial density profile of the simulated core (red points) compared to the analytic soliton solution (black line; fitted $r_{c}=2.32$ in code units). Both axes are shown in code units; the corresponding physical scaling for $m_{22}=2$ is given in Section IV (yielding $r_{c}\simeq 51$ pc and $M_{c}\simeq 7.8\times 10^{8}\,M_{\odot}$). The deviation at large radii reflects unrelaxed merger debris and finite-box/periodic-boundary effects, and does not affect the inner soliton core relevant for our analytic estimates.
@@ VI Discussion
VI.1 Resolution of the Lyman-$\alpha$ Tension
While a fiducial boson mass of $m_{22}\approx 2$ allows for a direct mapping where the black hole mass tracks the soliton mass ($M_{\rm BH}\approx M_{c}$), our analysis of the size–mass relation (Figure 4) reveals that the observed data are fully compatible with higher particle masses (e.g., $m_{22}\sim 20$5050, consistent with Lyman-$\alpha$ forest constraints; see, e.g., Iršič et al. 2017; Rogers & Peiris 2021) if LRDs represent an early progenitor phase where the central black hole has not yet consumed the entire soliton core. We note that the precise Lyman-$\alpha$ bound depends on assumptions about the IGM thermal history and on whether FDM constitutes all of the dark matter; mixed-DM scenarios with a subdominant FDM fraction can weaken the constraint (e.g., Kobayashi et al., 2017). In this scenario, the total soliton mass $M_{c}$ remains larger than the growing black hole mass $M_{\rm BH}$. By relaxing the strict assumption of $M_{\rm BH}=M_{c}$ and interpreting the horizontal scatter in Figure 4 (where the axis represents observed $M_{\rm BH}$) as a reflection of evolutionary stages ($M_{\rm BH}<M_{c}$), the parameter space naturally opens up. To clarify the apparent tension with Section II.1: the $m_{22}\approx 2$ baseline is strictly derived under a mature, direct-mapping assumption ($M_{\rm BH}\approx M_{c}$). However, during the formation stage, the host soliton mass can be substantially larger than the growing black hole ($M_{c}\gg M_{\rm BH}$). Therefore, the horizontal distribution of the observational data reflects black hole growth (increasing $M_{\rm BH}$) within larger, fixed soliton cores (fixed $m_{22}$), rather than varying $m_{22}$ at a fixed evolutionary stage. Under this progenitor interpretation, $m_{22}\gtrsim 20$ is fully consistent with both LRD sizes and Lyman-$\alpha$ limits, without contradicting the fiducial mature-phase estimate. The soliton provides the deep, compact potential well required to trigger the opacity crisis, while the observed black hole mass reflects only the collapsed fraction. Consequently, the apparent tension between the compact sizes of LRDs and the Lyman-$\alpha$ constraints may be substantially alleviated, suggesting that the FDM framework remains compatible with both the observed small-scale structure of LRDs and large-scale Lyman-$\alpha$ forest data.
@@ VI.2 Observational Signatures
  • Inverse Size–Mass Relation: For soliton cores at fixed $m_{22}$, the soliton scaling implies $r_{c}\propto M_{c}^{-1}$. Figure 4 illustrates this compactness floor and the observed population above it. Establishing an intrinsic inverse $r_{c}$$M_{c}$ relation as a discriminator will require independent constraints on host/core masses (beyond virial $M_{\rm BH}$ estimates) and careful modeling of how radiative transfer maps the intrinsic core scale to the observed $r_{e}$.
@@ VI.3 Demographics and Duty Cycle
A key question is whether the transient obscured phase implied by the “Opacity Crisis” can match the observed abundance of LRDs, $n_{\rm LRD}\sim 10^{-5}\,{\rm comoving\ Mpc^{-3}}$ (here cMpc denotes comoving Mpc). For massive halos at $z\sim 5$77, a representative number density is $n_{\rm halo}\sim 10^{-4}\,{\rm cMpc^{-3}}$ (e.g., Kokorev et al., 2024), suggesting an instantaneous occupancy fraction of order $f_{\rm occ}\equiv n_{\rm LRD}/n_{\rm halo}\sim 0.1$. More generally, the occupancy can be written as a time-averaged duty cycle,
@@
where $P_{\rm trig}$ is the probability that a given halo experiences an LRD-triggering event within the relevant redshift window, $N_{\rm evt}$ is the number of such episodes per halo, $t_{\rm evt}$ is the observable duration of each episode, and $t_{H}$ is the Hubble time at that epoch. While the characteristic dynamical time of the dense core can be short ($t_{\rm dyn}\sim\sqrt{r_{c}^{3}/(GM_{\rm enc})}\sim 10^{5}\,{\rm yr}$ for $r_{c}\sim 50$ pc and $M_{\rm enc}\sim 10^{9}\,M_{\odot}$), feedback, clumpy obscuration, and/or repeated episodes can extend the effective observable duration to $t_{\rm evt}\sim 10^{6}$$10^{7}\,{\rm yr}$. Matching $f_{\rm occ}\sim 0.1$ then requires $P_{\rm trig}N_{\rm evt}$ to be of order unity to several, motivating multiple merger-driven or inflow-driven obscured episodes during early assembly.
Figure 4: The size–mass plane of Little Red Dots (LRDs), comparing observed effective radii $r_{e}$ with illustrative FDM soliton compactness limits. The red shaded region indicates a “Soliton Formation Zone,” defined by the minimum radii implied by the soliton scaling for different boson masses. The horizontal axis represents the observed black hole mass ($M_{\rm BH}$), which does not necessarily equal the underlying soliton mass ($M_{c}$) during the early progenitor phase. The dashed red line shows our fiducial baseline ($m_{22}=2$) under the direct-mapping assumption ($M_{\rm BH}\approx M_{c}$). LRD measurements from Greene et al. (2024) and Kokorev et al. (2024) are overplotted. In the progenitor phase interpretation ($M_{\rm BH}<M_{c}$), the horizontal scatter of the observational data reflects the growth track of black holes at a fixed $m_{22}$ within larger, fixed soliton cores ($M_{c}\gg M_{\rm BH}$), rather than a varying $m_{22}$. This resolves the apparent tension with the $m_{22}\approx 2$ estimate in Section II.1, which strictly assumes a mature, direct-mapping ($M_{\rm BH}\approx M_{c}$) scenario. This interpretation allows the observed LRDs to be consistent with heavier boson masses ($m_{22}\gtrsim 20$, consistent with Lyman-$\alpha$ constraints) while still utilizing the deep soliton potential to trigger the initial collapse.
@@ VII Conclusion
We have presented a compact model in which LRD-like systems arise during a transient, heavily obscured phase inside FDM soliton cores. Combining size scaling with opacity constraints suggests $m_{22}\sim{\rm few}$. A conservative column estimate, including sensitivity analysis for baryon loading, indicates that reaching $N_{\rm H}\gtrsim 10^{24}\,{\rm cm^{-2}}$ typically implies radiative losses faster than dynamical support (or radiation-pressure dominance), disfavoring a long-lived static hot atmosphere and motivating rapid inflow or radiation-pressure-driven evolution. In progenitor-phase interpretations where $M_{\rm BH}<M_{c}$, the observed LRD population may also remain consistent with heavier boson masses ($m_{22}\gtrsim 20$), while still triggering the same opacity-crisis collapse mechanism.
Supplementary Notes: Detailed Derivations and Physical Interpretations
This document provides rigorous mathematical derivations for the scaling laws and physical timescales presented in the revised manuscript. Specifically, we explicitly derive the dynamical timescale based on the soliton gravitational potential and compare it with the radiative cooling timescale. We demonstrate that the inclusion of the soliton potential significantly deepens the central well compared to a naked black hole; however, even with this enhanced gravitational support, radiative losses dominate for massive cores, leading to the unavoidable “Opacity Crisis.”
Appendix A Soliton Scaling Laws (Section 2.1)
A.1 Equation 1: The Soliton Core Radius
Equation:
$r_{c}\approx 160\,\text{pc}\,m_{22}^{-2}\left(\frac{M_{s}}{10^{9}M_{\odot}}\right)^{-1}$ (A1)
Derivation: The Fuzzy Dark Matter (FDM) soliton represents the ground-state solution of the Schrödinger-Poisson system. The scale arises from the balance between quantum pressure (uncertainty principle) and self-gravity. The de Broglie wavelength is $\lambda_{dB}\sim\frac{\hbar}{mv}$. For a self-gravitating system, the virial velocity is $v\sim\sqrt{GM/r}$. Equating the quantum pressure scale $r\sim\lambda_{dB}$ with the gravitational scale:
$r\sim\frac{\hbar}{m\sqrt{GM/r}}\implies r^{1/2}\sim\frac{\hbar}{m\sqrt{GM}}\implies r\sim\frac{\hbar^{2}}{Gm^{2}M}$ (A2)
Substituting physical constants yields the inverse scaling $r_{c}\propto m^{-2}M_{s}^{-1}$.
Appendix B Timescale Hierarchy and Stability (Sections 2.2 & 2.4)
B.1 Dynamical Timescale: Soliton vs. Naked Black Hole
Definition: We define the dynamical time at the core radius $r_{c}$ based on the gravitational potential of the total enclosed mass. This distinguishes the “Soliton Container” scenario from a naked black hole scenario. Derivation: The dynamical time is defined as:
$t_{dyn}(r)\equiv\sqrt{\frac{r^{3}}{GM_{enc}(r)}}$ (B1)
Case A: Soliton-Dominated Potential (The LRD Scenario)
For a black hole embedded in a soliton, $M_{enc}(r_{c})\approx M_{BH}+M_{s}(<r_{c})$. In the progenitor phase or massive soliton limit where $M_{s}\gtrsim M_{BH}$, the potential is dominated by the soliton mass $M_{s}$. Substituting the soliton scaling relation $r_{c}\propto m^{-2}M_{s}^{-1}$:
$t_{dyn,sol}\propto\sqrt{\frac{(m^{-2}M_{s}^{-1})^{3}}{M_{s}}}=\sqrt{m^{-6}M_{s}^{-4}}=m^{-3}M_{s}^{-2}$ (B2)
Thus, for a fixed boson mass $m$, the dynamical support time scales as:
$t_{dyn}\propto M_{s}^{-2}$ (B3)
This indicates that more massive solitons have quadratically shorter dynamical times due to their deeper and more compact potentials.
Case B: Naked Black Hole (Comparison for Major Point 1)
In the absence of a soliton (a “naked” black hole), the enclosed mass at the same radius $r_{c}$ is simply $M_{enc}\approx M_{BH}$. Since $M_{enc,sol}>M_{enc,BH}$ (often significantly, as $M_{s}\sim 10^{9}M_{\odot}$ while $M_{BH}$ may be growing), we have:
$t_{dyn,sol}<t_{dyn,BH}$ (B4)
The soliton creates a deeper potential well, providing stronger gravitational support (shorter collapse time) than the black hole alone. The fact that the “Opacity Crisis” (instability) occurs even in the soliton case ($t_{cool}<t_{dyn,sol}$) implies that a static atmosphere is a fortiori impossible in the naked black hole case, where the dynamical support is weaker (longer $t_{dyn}$). The role of the soliton is therefore to define the compact physical scale ($r_{c}\sim 50$ pc) and density threshold where this crisis manifests.
B.2 Cooling Timescale Derivation
Gas Properties: We assume the gas initially traces the soliton potential in hydrostatic equilibrium.
  • Virial Temperature: $T\propto\frac{M_{s}}{r_{c}}\propto\frac{M_{s}}{M_{s}^{-1}}\propto M_{s}^{2}$.
  • Gas Density: Assuming gas mass $M_{g}\propto M_{s}$ fills the volume $r_{c}^{3}$:
    $n\approx\frac{M_{g}}{m_{p}r_{c}^{3}}\propto\frac{M_{s}}{(M_{s}^{-1})^{3}}\propto M_{s}^{4}$ (B5)
Cooling Time and Multiphase Scaling: The cooling timescale is given by $t_{\rm cool}\sim\frac{k_{B}T}{n\Lambda(T)}$. Given $n\propto M_{s}^{4}$ and $T\propto M_{s}^{2}$, the scaling depends on the cooling function $\Lambda(T)$:
  • Atomic Cooling Threshold ($T<10^{4}$ K): For $M_{s}<2.8\times 10^{7}M_{\odot}$, $\Lambda(T)$ is exponentially suppressed by the Boltzmann factor. Consequently, $t_{\rm cool}$ increases sharply as $M_{s}$ decreases, exceeding the dynamical time and creating a stable regime.
  • Line Cooling Regime ($10^{4}<T<10^{7}$ K): As implemented in our numerical model, $\Lambda(T)\propto T^{-0.6}$ in this interval. The scaling becomes:
    $t_{\rm cool}\propto\frac{T}{n\Lambda(T)}\propto\frac{T^{1.6}}{n}\propto\frac{(M_{s}^{2})^{1.6}}{M_{s}^{4}}=M_{s}^{-0.8}$ (B6)
    This explains the shallower slope of the cooling curve in the middle section of Figure 2.
  • Bremsstrahlung Regime ($T>10^{7}$ K): For massive cores ($M_{s}\sim 10^{9}M_{\odot}$), thermal bremsstrahlung dominates, so $\Lambda(T)\propto T^{1/2}$. Substituting the scalings yields:
    $t_{\rm cool}\propto\frac{T^{1/2}}{n}\propto\frac{(M_{s}^{2})^{1/2}}{M_{s}^{4}}=M_{s}^{-3}$ (B7)
    This confirms the steep slope $t_{\rm cool}\propto M_{s}^{-3}$ shown in Figure 2 for the high-mass regime.
B.3 The Crossover (The Opacity Crisis)
Comparing the slopes derived above against the dynamical support:
  • Dynamical Support: $t_{\rm dyn}\propto M_{s}^{-2}$
  • Radiative Cooling: $t_{\rm cool}\propto M_{s}^{-0.8}$ (intermediate) and $t_{\rm cool}\propto M_{s}^{-3}$ (high-mass)
The intersection $t_{\rm dyn}=t_{\rm cool}$ defines the stability limit. Numerically solving for our fiducial parameters ($m_{22}=2.0$), we find a critical mass scale:
$M_{\rm crit}\approx 2.8\times 10^{7}M_{\odot}$ (B8)
For $M_{s}>M_{\rm crit}$, the gas loses thermal energy faster than the potential can dynamically readjust ($t_{\rm cool}<t_{\rm dyn}$). This leads to catastrophic collapse rather than a static hot atmosphere, confirming the “Opacity Crisis”.
Appendix C Demographics (Section 6.3)
C.1 Equation 17: Occupancy Fraction
Equation:
$n_{LRD}\simeq n_{halo}P_{trig}N_{evt}\left(\frac{t_{evt}}{t_{H}}\right)$ (C1)
Derivation: This is a standard duty-cycle argument based on the Ergodic Hypothesis.
  • $n_{halo}$: Number density of potential host halos.
  • $P_{trig}$: Probability that a halo is triggered to host an LRD.
  • $N_{evt}$: Number of events per triggered halo.
  • $t_{evt}/t_{H}$: The fractional duration of the event relative to the Hubble time.
The observed number density is simply the total halo density weighted by the probability of catching a halo in the active, obscured LRD phase.